A biquadratic equation must have at least one real root. Is it true or false? give reasons.
1
Expert's answer
2020-03-02T09:23:17-0500

A biquadratic equation is a polynomial equation with degree 4 without having degree 1 and 3 terms. Since it is saying it "must have" at least one real root, it makes the statement false

To prove this I will provide a counter example.

Consider a biquadratic equation

"x^4+7x^2+12=0"

For finding roots , let "u=x^2"

Given equation will rewritten as

"u^2+7u+12=0"

"u^2+3u+4u+12=0"

"u(u+3)+4(u+3)=0"

"(u+3)(u+4)=0"

"i.e u=-3" or "u=-4"

"since" "u=x^2"

"x^2=-3" or "x^2=-4"

"x=±\sqrt{\smash[b]{-3}}" x="±\sqrt{\smash[b]{3}}i" or "x=±\sqrt{\smash[b]{-4}}" or "x=\u00b12i"


Which means a biquadratic equation can have all imaginary roots.